# TopoBDA: Towards Bezier Deformable Attention for Road Topology Understanding

### v2

@@ Abstract
Keywords:
Road Topology Understanding Centerline Detection Autonomous Driving Automated HDMap Generation 3D Lane Detection
@@ 1 Introduction
In autonomous driving, scenes primarily consist of two types of entities: dynamic and stationary. Dynamic entities encompass objects capable of movement and interaction, such as vehicles (cars, bicycles, motorcycles, trucks), and pedestrians. Stationary entities, on the other hand, are immobile objects and abstract structures such as lane markings, crosswalks, road signs, traffic lights, barriers, and road surfaces, that regulate traffic and ensure the safe and orderly movement of dynamic entities. Furthermore, abstract stationary entities -such as lanes and centerlines- are defined in relation to other stationary objects or specific rules of the driving environment.
@@
An alternative solution to this problem is the utilization of High-Definition Maps (HDMaps), which provide pre-computed maps of stationary entities. However, HDMaps are expensive to produce, cover only limited geographic areas, and cannot reflect recent changes on the road, requiring continuous updates. Furthermore, errors introduced by the Global Navigation Satellite System (GNSS) receiver on the vehicle can lead to localization inaccuracies. These inaccuracies may cause discrepancies between the actual position of the vehicle and the HDMap data, potentially resulting in drifts in the map-based guidance system. In response to these challenges, automatic HDMap construction has gained significant attention in recent years for two primary reasons [26, 39, 34, 3]. First, it can reduce reliance on HDMaps for autonomous driving. Second, it lowers the cost of creating and maintaining HDMaps, thereby minimizing the need for human effort.
@@
In the context of polyline structures, such as centerlines and lane dividers, the application of standard cross attention [3, 40], deformable cross attention [27, 66, 34, 46, 28, 39, 1, 82, 9, 42, 70, 76], and masked cross attention [53, 10, 82, 21] is common. Masked attention [8], as illustrated in Figure 1b, requires tuning hyperparameters for polyline width and uses a thresholding mechanism to differentiate between foreground and background. Moreover, to prevent the failure of the attention mechanism, a foreground check is performed for every query, increasing complexity and reducing deployment efficiency. In contrast, deformable attention [83] can be implemented in two primary ways. The first, Single-Point Deformable Attention (SPDA), shown in Figure 1c, limits attention for each polyline instance to a single point, typically the center point of the polyline’s bounding box [39, 66, 66]. However, this method restricts attention to a local scope, making it unsuitable for elongated, thin polyline structures.
The second method for implementing deformable attention, Multi-Point Deformable Attention (MPDA), as shown in Figure 1d, involves distributing attention around every predicted dense lane point [34, 77, 46, 70, 4]. From the MPDA perspective, there are two primary approaches: point query-based methods and instance query-based methods. Point query-based methods are inherently complex because each query represents a single point, and increasing the number of points per instance proportionally increases the complexity [34, 35, 46, 10, 25]. Conversely, instance query-based methods are more efficient [28, 77]. However, a gap in the literature exists, as instance query methods utilizing Bezier control points [66, 27] have not applied MPDA, relying instead on SPDA.
This study focuses on enhancing centerline detection performance within the broader context of road topology understanding. First, the integration of MPDA into Bezier keypoint-dependent transformer decoder structures is introduced. This adaptation is considered crucial because, unlike other dense polyline prediction methods, the Bezier representation [28] [77], leverages its inherently compact nature to improve the computational efficiency of polyline prediction. By incorporating MPDA into the Bezier keypoint-dependent transformer decoder, the method can more effectively handle elongated and thin polyline structures.
@@
In addition, consistent with the findings of the TopoMaskV2 study [21], the instance-mask formulation is employed to enhance the overall performance of road topology. However, unlike the direct approach used in TopoMaskV2, TopoBDA adopts an indirect application to reduce post-processing requirements. First, it has been demonstrated that incorporating the instance-mask formulation as an auxiliary loss significantly benefits the Bezier head. Second, during the Hungarian matching, using a Mask-L1 mix matcher [24], instead of a pure L1 matcher, has proven to be superior. These proposed mechanisms underscore the effectiveness of the indirect instance-mask formulation in improving road topology performance.
Sensor fusion has shown significant benefits in various domains, including 3D lane detection [44, 45] and HDMap element prediction [38, 78, 34, 39, 26]. Despite these advancements, there remains a notable gap in the literature regarding its application to road topology understanding. Previous studies have primarily focused on the performance gains of SDMap [43, 74], without exploring the potential of sensor fusion in this context. Our research is the first to investigate and comprehensively evaluate the effects of sensor fusion, utilizing both lidar and radar data for road topology understanding. Additionally, we analyze the benefits of integrating SDMap with lidar and camera sensors, highlighting the novel combination of lidar and SDMap, which has not been explored in the literature.
Auxiliary one-to-many set prediction loss strategy, adapted from hybrid matching techniques [20], is implemented for HDMap element prediction [35] and shown to improve convergence and performance without increasing the inference complexity. While, also employed in road topology understanding problem [66, 21], its quantitative benefits in this context have not been fully explored. Therefore, this research provides a comprehensive analysis of its impact on the proposed TopoBDA architecture.
With the inclusion of BDA, instance mask formulation, and one-to-many set prediction loss, TopoBDA achieves state-of-the-art results in the camera-only benchmark for both OpenLane-V1 and OpenLane-V2 datasets. Furthermore, when utilizing multi-modal data, TopoBDA attains state-of-the-art results in OpenLane-V2. To summarize, the key innovations and contributions of this work are as follows:
@@
  • Multi-Modal Fusion: Lidar and radar data are utilized for the first time specifically for road topology understanding. Additionally, the fusion of lidar with SDMap is analyzed, demonstrating the benefits of integrating multi-modal data. By fusing camera, lidar, and SDMap data, state-of-the-art results are achieved.
@@ 2.1.1 Perspective View Methods
Perspective view (PV) methods focus on detecting lane dividers from the PV and projecting them onto the ground using a homography matrix under the flat surface assumption. Despite the presence of various lane divider instances, semantic approaches are practical and effective, assuming a constant number of lane divider instances (e.g., 1st left, 2nd left, 1st right, 2nd right lane dividers). SCNN [50] adheres to this methodology and introduces a module that sequentially processes the rows and columns of the feature map. UFLD [54] transitions the formulation from pixel-based to grid-based, proposing row-based and column-based anchor formulations to enhance inference speed. LaneATT [56], inspired by object detection studies, devises an anchor concept specifically for lanes. The emergence of the CurveLanes dataset [68] has led to the adoption of instance-segmentation-based methods for the PV domain. CondLaneNet [36] introduces lane-specific methodologies such as offset prediction and row-wise formulation on top of the instance mask formulation in the PV domain. PolyLaneNet [57] and BezierLaneNet [11] employ polynomial and Bezier curve representations, respectively, to reduce post-processing efforts and improve curve learning.
@@
Recent studies have shifted towards directly predicting the 3D locations of lane dividers with the introduction of 3D lane divider datasets such as [72], OpenLane [5], and Apollo 3D Synthetic Lane [13]. This approach addresses the limitations of PV methods, specifically the flat world assumption due to the absence of depth information. Persformer [5] utilizes a deformable attention-based decoder with Inverse Perspective Mapping (IPM) and formulates a 3D anchor concept. CurveFormer [1] employs a sparse query design with a deformable attention mechanism and predicts polynomial parameters in the BEV domain. BEV-LaneDet [61] uses a keypoint concept with predicted offsets and groups the keypoints of the same lane instance with an embedding concept. PETRV2 [40] extends its sparse query design for lane detection. M2-3DLaneNet explores the benefits of integrating lidar and camera sensors. Another study [7] follows an instance-offset formulation in the BEV domain and aggregates offsets with a voting mechanism. LATR [46] introduces an end-to-end 3D lane detection framework that directly detects 3D lanes from front-view images using lane-aware queries and dynamic 3D ground positional embedding, significantly improving accuracy and efficiency over previous methods. GLane3D [84] utilizes a graph-based approach with keypoints and directed connections, achieving enhanced cross-dataset generalization performance.
@@ 2.2 HDMap Element Prediction
The prediction of High-Definition Map (HDMap) elements, such as lane dividers, road dividers, and pedestrian crossings, is essential for autonomous driving. Various methods have been developed to enhance the accuracy and efficiency of HDMap element prediction. HDMapNet [26] transforms Bird’s Eye View (BEV) semantic segmentation into BEV instance segmentation through extensive post-processing, leveraging predicted instance embeddings and directional information. VectorMapNet [39] introduces an end-to-end vectorized HD map learning pipeline that predicts sparse polylines from sensor observations, explicitly modeling spatial relationships between map elements, thus eliminating the need for dense rasterized segmentation and heuristic post-processing. Contrary to the autoregressive structure of VectorMapNet, MapTR [34] directly predicts points on polylines or polygons using a permutation-invariant Hungarian matcher, combining point query and instance query concepts, achieving real-time inference speeds and robust performance in complex driving scenes. InstaGraM [55] redefines polyline detection as a graph problem, where keypoints are vertices and their connections are edges, leveraging graph neural networks to enhance accuracy and robustness. MGMap [38] uses instance masks to generate map element queries and refine features with mask outputs, significantly improving performance over baseline methods. MapVR [78] rasterizes MapTR outputs and applies instance segmentation loss to address keypoint-based method limitations, enhancing performance without extra computational cost during inference. ADMap [16] employs instance interactive attention and vector direction difference loss to reduce point sequence jitter, enhancing map accuracy and stability. BeMapNet [53] models map elements as multiple piecewise curves using Bezier curves, eliminating the need for post-processing and achieving superior performance on existing benchmarks. StreamMapNet [77] explores the temporal aspects of HDMap element prediction, using temporal information as propagated instance queries and warped BEV features in a recurrent manner. MapTracker [4] formulates mapping as a tracking task, maintaining multiple memory latents to ensure consistent reconstructions over time, significantly outperforming existing methods on consistency-aware metrics. Recent advancements include PriorMapNet [60], which enhances online vectorized HD map construction by incorporating priors, and HIMap [82], which integrates point-level and element-level information to improve prediction accuracy. Additionally, the Multi-Session High-Definition Map-Monitoring System [64] employs machine learning algorithms to track and update map elements across multiple sessions, ensuring HD maps remain accurate and up-to-date. The Ultra-fast Semantic Map Perception [71] leverages both camera and LiDAR data to achieve real-time performance, featuring an orthogonal projection subspace for fast semantic segmentation and a Bayesian framework for enhanced global semantic fusion.
@@ 2.3.1 Centerline Concept
STSU [3] introduced a novel approach for extracting a directed graph of the local road network in bird’s-eye-view (BEV) coordinates from a single onboard camera image, significantly improving traffic scene understanding. CenterLineDet [69] uses a transformer network to detect lane centerlines with vehicle-mounted sensors, effectively handling complex graph topologies such as lane intersections. LaneGAP [33] models lane graphs path-wise, preserving lane continuity and improving the accuracy of lane graph construction. MapTRV2 [35] enhances centerline prediction efficiency by treating lane centerlines as paths and incorporating semantic-aware shape modeling. SMERF [43] integrates Standard Definition (SD) maps by tokenizing map elements using a transformer encoder, and employs these tokens in the cross-attention mechanism of a transformer decoder to improve lane detection and topology prediction. In contrast, TopoSD [74] uses both map tokens and feature maps extracted from SD maps to enrich BEV features. SMART [75] uniquely leverages SD and satellite maps to learn robust map priors, enhancing lane topology reasoning for autonomous driving without relying on consistent sensor configurations. LaneSegNet [28] introduces the concept of lane segments, combining geometry and topology information to provide a comprehensive representation of road structures.

### v3

@@ Abstract
Project page: https://artest08.github.io/TopoBDA.github.io/
@@ 1 Introduction
In autonomous driving, scenes primarily consist of two types of entities: dynamic and stationary. Dynamic entities encompass objects capable of movement and interaction, such as vehicles (cars, bicycles, motorcycles, trucks), and pedestrians. Stationary entities, on the other hand, are immobile objects and abstract structures such as lane markings, crosswalks, road signs, traffic lights, barriers, and road surfaces that regulate traffic and ensure the safe and orderly movement of dynamic entities. Furthermore, abstract stationary entities -such as lanes and centerlines- are defined in relation to other stationary objects or specific rules of the driving environment.
@@
An alternative solution to this problem is the utilization of High-Definition Maps (HDMaps), which provide pre-computed maps of stationary entities. However, HDMaps are expensive to produce, cover only limited geographic areas, and cannot reflect recent changes on the road, requiring continuous updates. Furthermore, errors introduced by the Global Navigation Satellite System (GNSS) receiver on the vehicle can lead to localization inaccuracies. These inaccuracies may cause discrepancies between the actual position of the vehicle and the HDMap data, potentially resulting in drifts in the map-based guidance system. In response to these challenges, automatic HDMap construction has gained significant attention in recent years for two primary reasons [li_hdmapnet_2022, liu_vectormapnet_2023, liao_maptr_2023, can_structured_2021]. First, it can reduce reliance on HDMaps for autonomous driving. Second, it lowers the cost of creating and maintaining HDMaps, thereby minimizing the need for human effort.
@@
In the context of polyline structures, such as centerlines and lane dividers, the application of standard cross attention [can_structured_2021, liu_petrv2_2023], deformable cross attention [li_graph-based_2023, wu_topomlp_2024, liao_maptr_2023, luo_latr_2023, li_lanesegnet_2024, liu_vectormapnet_2023, bai_curveformer_2023, zhou_himap_2024, choi_mask2map_2024, liu_leveraging_2024, xu_insmapper_2024, yu_scalablemap_2023], and masked cross attention [qiao_end--end_2023, ding_pivotnet_2023, zhou_himap_2024, kalfaoglu_topomaskv2_2024] is common. Masked attention [cheng_masked-attention_2022], as illustrated in Figure 1b, requires tuning hyperparameters for polyline width and uses a thresholding mechanism to differentiate between foreground and background. Moreover, to prevent the failure of the attention mechanism, a foreground check is performed for every query, increasing complexity and reducing deployment efficiency. In contrast, deformable attention [zhu_deformable_2020] can be implemented in two primary ways. The first, Single-Point Deformable Attention (SPDA), shown in Figure 1c, limits attention for each polyline instance to a single point, typically a learnable query embedding or the center point of the polyline’s bounding box [liu_vectormapnet_2023, li_graph-based_2023, wu_topomlp_2024]. However, this method restricts attention to a local scope, making it unsuitable for elongated, thin polyline structures.
The second method for implementing deformable attention, Multi-Point Deformable Attention (MPDA), as shown in Figure 1d, involves distributing attention around every predicted dense lane point [liao_maptr_2023, yuan_streammapnet_2024, luo_latr_2023, xu_insmapper_2024, chen_maptracker_2025]. From the MPDA perspective, there are two primary approaches: point query-based methods and instance query-based methods. Point query-based methods are inherently complex because each query represents a single point, and increasing the number of points per instance proportionally increases the complexity [liao_maptr_2023, liao_maptrv2_2024, luo_latr_2023, ding_pivotnet_2023, li_enhancing_2024]. Conversely, instance query-based methods are more efficient [li_graph-based_2023, li_lanesegnet_2024, yuan_streammapnet_2024]. However, a gap in the literature exists, as instance query methods utilizing Bezier control points [wu_topomlp_2024] have not applied MPDA, relying instead on SPDA.
This study focuses on enhancing centerline detection performance within the broader context of road topology understanding. First, the integration of MPDA into Bezier keypoint-dependent transformer decoder structures is introduced. This adaptation is considered crucial because, unlike other dense polyline prediction methods [li_graph-based_2023, li_lanesegnet_2024, yuan_streammapnet_2024], the Bezier representation leverages its inherently compact nature to improve the computational efficiency of polyline prediction. By incorporating MPDA into the Bezier keypoint-dependent transformer decoder, the method can more effectively handle elongated and thin polyline structures.
@@
In addition, consistent with the findings of the TopoMaskV2 study [kalfaoglu_topomaskv2_2024], the instance-mask formulation is employed to enhance the overall performance of road topology. However, unlike the direct approach used in TopoMaskV2, TopoBDA adopts an indirect application to reduce post-processing requirements. First, it has been demonstrated that incorporating the instance-mask formulation as an auxiliary loss significantly benefits the Bezier head. Second, during the Hungarian matching, using a Mask-L1 mix matcher [li_mask_2023], instead of a pure L1 matcher, has proven to be superior. These proposed mechanisms underscore the effectiveness of the indirect instance-mask formulation in improving road topology performance.
Sensor fusion has shown significant benefits in various domains, including 3D lane detection [luo_dv-3dlane_2024, luo_m2-3dlanenet_2022] and HDMap element prediction [liu_mgmap_2024, zhang_online_2024, liao_maptr_2023, liu_vectormapnet_2023, li_hdmapnet_2022]. Despite these advancements, there remains a notable gap in the literature regarding its application to road topology understanding. Previous studies have primarily focused on the performance gains of SDMap [luo_augmenting_2023, yang_toposd_2024], without exploring the potential of sensor fusion in this context. Our research is the first to investigate and comprehensively evaluate the effects of sensor fusion, utilizing both lidar and radar data for road topology understanding. Additionally, we analyze the benefits of integrating SDMap with lidar and camera sensors, highlighting the novel combination of lidar and SDMap, which has not been explored in the literature.
Auxiliary one-to-many set prediction loss strategy, adapted from hybrid matching techniques [jia_detrs_2023], is implemented for HDMap element prediction [liao_maptrv2_2024] and shown to improve convergence and performance without increasing the inference complexity. While also employed in road topology understanding problem [wu_topomlp_2024, kalfaoglu_topomaskv2_2024], its quantitative benefits in this context have not been fully explored. Therefore, this research provides a comprehensive analysis of its impact on the proposed TopoBDA architecture.
With the inclusion of BDA, instance mask formulation, and one-to-many set prediction loss, TopoBDA achieves state-of-the-art results in the camera-only benchmark for both OpenLane-V1 and OpenLane-V2 datasets. Furthermore, when utilizing multi-modal data, TopoBDA attains state-of-the-art results in OpenLane-V2. The contribution of this study is detailed in Supplementary Section S.8. To summarize, the key innovations and contributions of this work are as follows:
@@
  • Multi-Modal Fusion: Lidar and radar data are utilized for the first time, specifically for road topology understanding. Additionally, the fusion of lidar with SDMap is analyzed, demonstrating the benefits of integrating multi-modal data. By fusing camera, lidar, and SDMap data, state-of-the-art results are achieved.
@@ 2.1.1 Perspective View Methods
Perspective view (PV) methods focus on detecting lane dividers from the PV and projecting them onto the ground using a homography matrix under the flat surface assumption. Despite the presence of various lane divider instances, semantic approaches are practical and effective, assuming a constant number of lane divider instances (e.g., 1st left, 2nd left, 1st right, 2nd right lane dividers). SCNN [pan_spatial_2018] adheres to this methodology and introduces a module that sequentially processes the rows and columns of the feature map. UFLD [qin_ultra_2022] transitions the formulation from pixel-based to grid-based, proposing row-based and column-based anchor formulations to enhance inference speed. LaneATT [tabelini_keep_2021], inspired by object detection studies, devises an anchor concept specifically for lanes. The emergence of the CurveLanes dataset [xu_curvelane-nas_2020] has led to the adoption of instance-segmentation-based methods for the PV domain. CondLaneNet [liu_condlanenet_2021] introduces lane-specific methodologies such as offset prediction and row-wise formulation on top of the instance mask formulation in the PV domain. PolyLaneNet [tabelini_polylanenet_2021] and BezierLaneNet [feng_rethinking_2022] employ polynomial and Bezier curve representations, respectively, to reduce post-processing efforts and improve curve learning. BezierFormer [dong_bezierformer_2024] introduces a Bezier curve attention mechanism that aggregates sampled features along the curve across all attention heads in front-view image space, whereas our TopoBDA adopts a head-specific reference point design in bird’s-eye view, enabling more efficient feature extraction in a multi-camera setup.
@@
Recent studies have shifted towards directly predicting the 3D locations of lane dividers with the introduction of 3D lane divider datasets such as [yan_once-3dlanes_2022], OpenLane [chen_persformer_2022], and Apollo 3D Synthetic Lane [guo_gen-lanenet_2020]. This approach addresses the limitations of PV methods, specifically the flat world assumption due to the absence of depth information. Persformer [chen_persformer_2022] utilizes a deformable attention-based decoder with Inverse Perspective Mapping (IPM) and formulates a 3D anchor concept. CurveFormer [bai_curveformer_2023] employs a sparse query design with a deformable attention mechanism and predicts polynomial parameters in the BEV domain. BEV-LaneDet [wang_bev-lanedet_2023] uses a keypoint concept with predicted offsets and groups the keypoints of the same lane instance with an embedding concept. PETRV2 [liu_petrv2_2023] extends its sparse query design for lane detection. M2-3DLaneNet explores the benefits of integrating lidar and camera sensors. Another study [chen_efficient_2023] follows an instance-offset formulation in the BEV domain and aggregates offsets with a voting mechanism. LATR [luo_latr_2023] introduces an end-to-end 3D lane detection framework that directly detects 3D lanes from front-view images using lane-aware queries and dynamic 3D ground positional embedding, significantly improving accuracy and efficiency over previous methods. GLane3D [ozturk_glane3d_2025] utilizes a graph-based approach with keypoints and directed connections, achieving enhanced cross-dataset generalization performance.
@@ 2.2 HDMap Element Prediction
The prediction of High-Definition Map (HDMap) elements, such as lane dividers, road dividers, and pedestrian crossings, is essential for autonomous driving. Various methods have been developed to enhance the accuracy and efficiency of HDMap element prediction. HDMapNet [li_hdmapnet_2022] transforms Bird’s Eye View (BEV) semantic segmentation into BEV instance segmentation through extensive post-processing, leveraging predicted instance embeddings and directional information. VectorMapNet [liu_vectormapnet_2023] introduces an end-to-end vectorized HD map learning pipeline that predicts sparse polylines from sensor observations, explicitly modeling spatial relationships between map elements, thus eliminating the need for dense rasterized segmentation and heuristic post-processing. Contrary to the autoregressive structure of VectorMapNet, MapTR [liao_maptr_2023] directly predicts points on polylines or polygons using a permutation-invariant Hungarian matcher, combining point query and instance query concepts, achieving real-time inference speeds and robust performance in complex driving scenes. InstaGraM [shin_instagram_2023] redefines polyline detection as a graph problem, where keypoints are vertices and their connections are edges, leveraging graph neural networks to enhance accuracy and robustness. MGMap [liu_mgmap_2024] uses instance masks to generate map element queries and refine features with mask outputs, significantly improving performance over baseline methods. MapVR [zhang_online_2024] rasterizes MapTR outputs and applies instance segmentation loss to address keypoint-based method limitations, enhancing performance without extra computational cost during inference. ADMap [hu_admap_2024] employs instance interactive attention and vector direction difference loss to reduce point sequence jitter, enhancing map accuracy and stability. BeMapNet [qiao_end--end_2023] models map elements as multiple piecewise curves using Bezier curves, eliminating the need for post-processing and achieving superior performance on existing benchmarks. StreamMapNet [yuan_streammapnet_2024] explores the temporal aspects of HDMap element prediction, using temporal information as propagated instance queries and warped BEV features in a recurrent manner. MapTracker [chen_maptracker_2025] formulates mapping as a tracking task, maintaining multiple memory latents to ensure consistent reconstructions over time, significantly outperforming existing methods on consistency-aware metrics. Recent advancements include PriorMapNet [wang_priormapnet_2024], which enhances online vectorized HD map construction by incorporating priors, and HIMap [zhou_himap_2024], which integrates point-level and element-level information to improve prediction accuracy. Additionally, the Multi-Session High-Definition Map-Monitoring System [wijaya_multi-session_2023] employs machine learning algorithms to track and update map elements across multiple sessions, ensuring HD maps remain accurate and up-to-date. The Ultra-fast Semantic Map Perception [xu_ultra-fast_2024] leverages both camera and LiDAR data to achieve real-time performance, featuring an orthogonal projection subspace for fast semantic segmentation and a Bayesian framework for enhanced global semantic fusion.
@@ 2.3.1 Centerline Concept
STSU [can_structured_2021] introduced a novel approach for extracting a directed graph of the local road network in bird’s-eye-view (BEV) coordinates from a single onboard camera image, significantly improving traffic scene understanding. CenterLineDet [xu_centerlinedet_2022] uses a transformer network to detect lane centerlines with vehicle-mounted sensors, effectively handling complex graph topologies such as lane intersections. LaneGAP [liao_lane_2023] models lane graphs path-wise, preserving lane continuity and improving the accuracy of lane graph construction. MapTRV2 [liao_maptrv2_2024] enhances centerline prediction efficiency by treating lane centerlines as paths and incorporating semantic-aware shape modeling. SMERF [luo_augmenting_2023] integrates Standard Definition (SD) maps by tokenizing map elements using a transformer encoder, and employs these tokens in the cross-attention mechanism of a transformer decoder to improve lane detection and topology prediction. In contrast, TopoSD [yang_toposd_2024] uses both map tokens and feature maps extracted from SD maps to enrich BEV features. SMART [ye_smart_2025] uniquely leverages SD and satellite maps to learn robust map priors, enhancing lane topology reasoning for autonomous driving without relying on consistent sensor configurations. LaneSegNet [li_lanesegnet_2024] introduces the concept of lane segments, combining geometry and topology information to provide a comprehensive representation of road structures.